### 13 Mar 2026 v1

@@ 3.2 Mechanism Analysis: Explaining TabPFN’s active-learning performance
In terms of uncertainty quantification—which governs the “exploration” efficiency of the active learning loop—we utilized the NLL, Prediction Interval Coverage Probability (PICP), Mean Prediction Interval Width (MPIW), and the Area Under the Sparsification Error (AUSE) curve. PICP measures the probability that the true value falls within the predicted confidence interval (target 0.95), while MPIW measures the sharpness of those intervals. A useful uncertainty estimate must balance high coverage with narrow width. First we found that TabPFN achieved the lowest NLL (2.34) compared to the Gaussian Process (3.46) and Random Forest (3.11), a result that is significant because NLL is a proper scoring rule that penalizes both prediction error and poor uncertainty calibration; this confirms that TabPFN’s predicted posterior distributions are the most faithful to the true data distribution. Next, while the Gaussian Process achieved high coverage (PICP 0.93), it did so by producing excessively wide intervals (MPIW 32.50), rendering the uncertainty signal uninformative for distinguishing between candidates. In contrast, TabPFN achieved comparable coverage with 3x sharper intervals (MPIW 11.16) and the lowest AUSE (0.83) compared to 15.58 of GP and 1.77 of RF. The low AUSE score is particularly telling, as it quantifies the correlation between estimated uncertainty and actual prediction error; a low score implies that the model is “honestly” uncertain only when it is likely to be wrong. This superior calibration allows the TabPFN-driven acquisition function to efficiently navigate the chemical space, avoiding the “blind” exploration often caused by the over-conservative uncertainty estimates of standard Gaussian Processes.
@@
We also evaluated the peformances of TabPFN and the baseline ML models over the more challenging material design for higher LTC using the concentration features as shown in Table 6, which presents five-fold cross-validation results evaluating both predictive accuracy and uncertainty quantification quality for the three surrogate models on the LTC dataset with concentration features. While all three models achieve comparable predictive accuracy — TabPFN and RF both attain RMSE 48.7 and R2 = 0.18, modestly outperforming GP whose R2 of 0.01 indicates near-zero predictive power — the critical differentiator lies in the quality of uncertainty estimation, which directly governs active learning query selection. GP’s near-perfect PICP of 0.99 is deceptive: it is achieved trivially through extremely wide prediction intervals (MPIW = 237.23, nearly 10 larger than TabPFN’s 33.16), meaning GP’s uncertainty covers almost the entire output range regardless of input and provides no meaningful signal for distinguishing informative from uninformative candidates. This pathological overestimation of uncertainty explains GP’s poor active learning performance on ltc_conc in Table 4, where it requires 56.55 extra evaluations on average — 29.4% more than TabPFN. RF achieves a tighter MPIW of 24.37 but significantly undercoveres at PICP = 0.86, falling well short of the nominal 95% level and indicating that its heuristic ensemble variance systematically underestimates true predictive uncertainty, which in turn causes the acquisition function to over-exploit overconfident predictions and miss the global optimum — consistent with RF requiring 61.8 extra evaluations in Table 4. TabPFN achieves the best balance: a PICP of 0.94 closest to the nominal 95% coverage, a well-controlled MPIW of 33.16, the lowest NLL of 3.22 (less than half of GP’s 5.81 and less than half of RF’s 7.53), and the lowest AUSE of 2.16 — less than half of RF’s 4.50 and five times lower than GP’s 10.80. The AUSE metric is particularly relevant here as it directly measures how well-ordered the uncertainty estimates are for active learning: a low AUSE means that samples flagged as uncertain by the surrogate are indeed the ones with the largest prediction errors, enabling the acquisition function to query the most informative compositions efficiently. This superior uncertainty calibration is the mechanistic explanation for TabPFN’s 29.4% and 35.4% savings in extra evaluations relative to GP and RF respectively on ltc_conc in Table 4, and underscores that predictive accuracy alone is insufficient to characterize surrogate quality for active learning — reliable and calibrated uncertainty quantification is equally, if not more, important.
@@ 4 Discussion and Conclusion
ICAL is also well-positioned for integration into self-driving laboratory (SDL) platforms [45, 9, 42, 1], where the surrogate model must make reliable query decisions in real time with continuously growing but always small labeled datasets. TabPFN’s zero-retraining inference — performing a complete Bayesian update in a single forward pass as new experimental results are added to the context — aligns naturally with the closed-loop, iterative experimentation paradigm of SDLs, eliminating the retraining latency that can become a bottleneck in high-throughput autonomous workflows. Furthermore, ICAL can be straightforwardly extended from pool-based screening to open-ended generative materials design [12, 30], where candidate materials are proposed by a generative model and TabPFN serves as the screening surrogate to prioritize which generated candidates are worth evaluating experimentally.
Several limitations of the current work point toward promising directions for future research. The performance reversal on the ltc_structure dataset, where RF outperforms TabPFN, suggests that structural features encode an inductive bias that TabPFN’s tabular meta-prior does not currently capture. Extending TabPFN’s pre-training distribution to include structure-aware representations, or developing hybrid architectures that combine TabPFN’s probabilistic inference with graph neural network encoders for structural features, could address this limitation. The identified minimum data threshold effect also motivates the development of adaptive initialization strategies that ensure sufficient compositional coverage before the AL loop begins. Finally, while we evaluate single-objective AL in this work, extending ICAL to multi-objective settings [19] — where practitioners simultaneously optimize multiple competing properties such as strength and ductility, or conductivity and corrosion resistance — represents a natural and practically important extension that we leave for future work.

### 17 Mar 2026 v2

@@ 3.2 Mechanism Analysis: Explaining TabPFN’s active-learning performance
In terms of uncertainty quantification, which governs the “exploration” efficiency of the active learning loop, we utilized the NLL, Prediction Interval Coverage Probability (PICP), Mean Prediction Interval Width (MPIW), and the Area Under the Sparsification Error (AUSE) curve. PICP measures the probability that the true value falls within the predicted confidence interval (target 0.95), while MPIW measures the sharpness of those intervals. A useful uncertainty estimate must balance high coverage with narrow width. First we found that TabPFN achieved the lowest NLL (2.34) compared to the Gaussian Process (3.46) and Random Forest (3.11), a result that is significant because NLL is a proper scoring rule that penalizes both prediction error and poor uncertainty calibration; this confirms that TabPFN’s predicted posterior distributions are the most faithful to the true data distribution. Next, while the Gaussian Process achieved high coverage (PICP 0.93), it did so by producing excessively wide intervals (MPIW 32.50), rendering the uncertainty signal uninformative for distinguishing between candidates. In contrast, TabPFN achieved comparable coverage with 3x sharper intervals (MPIW 11.16) and the lowest AUSE (0.83) compared to 15.58 of GP and 1.77 of RF. The low AUSE score is particularly telling, as it quantifies the correlation between estimated uncertainty and actual prediction error; a low score implies that the model is “honestly” uncertain only when it is likely to be wrong. This superior calibration allows the TabPFN-driven acquisition function to efficiently navigate the chemical space, avoiding the “blind” exploration often caused by the over-conservative uncertainty estimates of standard Gaussian Processes.
@@
We also evaluated the peformances of TabPFN and the baseline ML models over the more challenging material design for higher LTC using the concentration features as shown in Table 6, which presents five-fold cross-validation results evaluating both predictive accuracy and uncertainty quantification quality for the three surrogate models on the LTC dataset with concentration features. While all three models achieve comparable predictive accuracy (TabPFN and RF both attain RMSE 48.7 and R2 = 0.18, modestly outperforming GP whose R2 of 0.01 indicates near-zero predictive power), the critical differentiator lies in the quality of uncertainty estimation, which directly governs active learning query selection. GP’s near-perfect PICP of 0.99 is deceptive: it is achieved trivially through extremely wide prediction intervals (MPIW = 237.23, nearly 10 larger than TabPFN’s 33.16), meaning GP’s uncertainty covers almost the entire output range regardless of input and provides no meaningful signal for distinguishing informative from uninformative candidates. This pathological overestimation of uncertainty explains GP’s poor active learning performance on ltc_conc in Table 4, where it requires 56.55 extra evaluations on average — 29.4% more than TabPFN. RF achieves a tighter MPIW of 24.37 but significantly undercoveres at PICP = 0.86, falling well short of the nominal 95% level and indicating that its heuristic ensemble variance systematically underestimates true predictive uncertainty, which in turn causes the acquisition function to over-exploit overconfident predictions and miss the global optimum , consistenting with RF requiring 61.8 extra evaluations in Table 4. TabPFN achieves the best balance: a PICP of 0.94 closest to the nominal 95% coverage, a well-controlled MPIW of 33.16, the lowest NLL of 3.22 (less than half of GP’s 5.81 and less than half of RF’s 7.53), and the lowest AUSE of 2.16 — less than half of RF’s 4.50 and five times lower than GP’s 10.80. The AUSE metric is particularly relevant here as it directly measures how well-ordered the uncertainty estimates are for active learning: a low AUSE means that samples flagged as uncertain by the surrogate are indeed the ones with the largest prediction errors, enabling the acquisition function to query the most informative compositions efficiently. This superior uncertainty calibration is the mechanistic explanation for TabPFN’s 29.4% and 35.4% savings in extra evaluations relative to GP and RF respectively on ltc_conc in Table 4, and underscores that predictive accuracy alone is insufficient to characterize surrogate quality for active learning because reliable and calibrated uncertainty quantification is equally, if not more, important.
@@ 4 Discussion and Conclusion
These results suggest that foundation models can address the small-data surrogate modeling challenge in experimental discovery.
@@
ICAL is also well-positioned for integration into self-driving laboratory (SDL) platforms [45, 9, 42, 1], where the surrogate model must make reliable query decisions in real time with continuously growing but always small labeled datasets. TabPFN’s zero-retraining inference can perform a complete Bayesian update in a single forward pass as new experimental results are added to the context, which aligns naturally with the closed-loop, iterative experimentation paradigm of SDLs, eliminating the retraining latency that can become a bottleneck in high-throughput autonomous workflows. Furthermore, ICAL can be straightforwardly extended from pool-based screening to open-ended generative materials design [12, 30], where candidate materials are proposed by a generative model and TabPFN serves as the screening surrogate to prioritize which generated candidates are worth evaluating experimentally.
Several limitations of the current work point toward promising directions for future research. The performance reversal on the ltc_structure dataset, where RF outperforms TabPFN, suggests that structural features encode an inductive bias that TabPFN’s tabular meta-prior does not currently capture. Extending TabPFN’s pre-training distribution to include structure-aware representations, or developing hybrid architectures that combine TabPFN’s probabilistic inference with graph neural network encoders for structural features, could address this limitation. The identified minimum data threshold effect also motivates the development of adaptive initialization strategies that ensure sufficient compositional coverage before the AL loop begins. Finally, while we evaluate single-objective AL in this work, extending ICAL to multi-objective settings [19] represents a natural and practically important extension that we leave for future work, where practitioners simultaneously optimize multiple competing properties such as strength and ductility, or conductivity and corrosion resistance.